### v5

@@ 3 Methodology
We detail wintermute, our prototype to rapidly evaluate privilege-escalation capabilities of LLMs. We are using a Linux Privilege Escalation benchmark suite detailed in Section 4 that provides these security guarantees, and its test scenarios are detailed in Table 3. In addition to our specialized Linux Priv-Esc research area, other benchmarks are not publicly available for local testing or do not provide fine-grained vulnerabilities to better analyze LLM behavior.
3.1 Benchmark Workflow
Our prototype allows for fully automated evaluation of an LLM’s privilege-escalation capabilities. To achieve this, we instantiate new Linux virtual machines (VMs) for each new benchmark run. Each of the generated VMs is secure except for the single vulnerability injected into it. The virtual machines are subsequently used as targets for the configured LLM and privilege attacks are performed as detailed in Section 3.2. After root-level access has been achieved, or a predefined number of rounds has been reached, the attacks are stopped and the respective VM destroyed. We keep the log information according to Section 3.3 for later analysis.
@@ 3.1 Benchmark Workflow
3.2 Autonomous Hacking Agent
Figure 1: Relationship between prompts and stored data. Blue boxes are LLM-callouts, orange boxes denote python modules.
Figure 2: Relationship between prompts and stored data.
@@ 3.2 Autonomous Hacking Agent
Figure 3: Prompt used to query next command.
@@
Figure 4: Prompt used to update the state/worldview.
hackingBuddyGPT is a Python program that supervises and controls the privilege-escalation attempts. It connects to the target VM through SSH as well as to the used LLMs through an OpenAI-compatible HTTP API. It is also responsible for collecting and storing log information for subsequent analysis.
All decision-making is performed by LLMs through prompting. The main prompt, next-cmd, detailed in Figure 3, includes information currently known by the agent and asks for the next command to be executed on the target machine. The prompts are implemented as Python Mako templates where context information and the current agents worldview are injected through text variables. The interactions between the prompts and the stored data are shown in Figure 1.
Figure 1 also includes the variable capabilities contains a list of available capabilities, i.e., the actions that an LLM can request to interact with the target system. Our prototype provides two actions: execute_command and test_credentials. When using the former, the LLM provides the command to be executed on the target system as parameters, e.g., “execute_cmd ls”. With the latter, the LLM can test the validity of credentials on the target system, e.g., “test_credentials root password”.
The agent’s world view is represented through the variables history, state and guidance. Each of these optional variables can be configured through configuration options. History contains the output of all prior executed commands, i.e., resembles the shell history as seen within an terminal emulator. It is automatically truncated to fit the used LLM’s context size.
The variable state is created using the optional second prompt “update-state” shown in Figure 4. Initially, the state is an empty list. After each executed system command, the LLM is shown the current state, the executed command and its output. It is then tasked with generating a new state, i.e., an updated state list containing information that it has gathered about the target system. If state is used instead of history, the used context size should thus be minimized at the cost of performing an additional LLM prompt as now both next-cmd and update-state are called for each round. This should be suitable for long-running agents where otherwise the size of the accumulated history would incur substantial costs. This implements a Reflection or Iterated Amplification pattern.
Finally, guidance can be used to prime the LLM towards a designated area or vulnerability class. This was implemented to emulate the high-level guidance typically provided to the Executor module in Planner architectures.
@@ 3.3 Collected Metrics
LLM query-specific data contains the type of query, the executed LLM prompt and its answer, the cost of performing the LLM prompt measured in elapsed time and utilized token counts, as well as the capability to be executed against the target system and its resulting response. For example, the captured data for command next-cmd stores the LLM’s prompt and answer, but also the executed capability and its result.
@@
The collected data allow us to perform both quantitative analysis, e.g., number of rounds needed for priv-esc, as well as qualitative analysis, e.g., quality of the LLM-derived system commands. As cloud-based models are typically priced by utilized tokens, capturing those allows us to analyze potential costs of LLM-guided penetration testing.
3.4 Experiment Design
Our experiments were designed with multiple goals in mind: a) to compare the efficacy of different LLM families for penetration testing, b) identify test scenarios problematic for LLMs, c) to evaluate the impact of memory and guidance upon model performance, d) evaluate the impact of context size differences upon performance, and e) the feasibility of small language models for penetration-testing.
State vs. History. Our baseline collects shell history until the respective LLM context is filled, inflicting substantial costs over time. We will analyze differences in the fill rate differential between different model families. An alternative is to utilize a compressed state by asking the LLM to summarize its current worldview based upon its prior worldview and the result of the last executed system command. As the summarized state grows slower than the original history, this effectively reduces context size at the cost of an additional LLM prompt call per round.
Impact of High-Level Guidance. The potential action state for LLMs driving Linux privilege-escalation is immense, creating the peril of LLMs not covering potential attack vectors. Our previous research indicates that providing high-level guidance substantially improves LLM performance (Happe and Cito, 2023a). To verify this, we implement two high-level guidance mechanisms utilizing the guidance prompt variable: on one hand, we will utilize hints provided by the used benchmark. For each scenario, a single hint is provided, e.g., for the suid scenario the hint is “there might be some exploitable suid binary on the system”. On the other hand, we implemented automated guidance based upon system enumeration. To implement this, the agent initially executes a system enumeration script (the mentioned linux-smart-enumeration.sh) and subsequently uses an LLM prompt to aggregate that enumeration information into up to three attack vector hints for hacking the respective system. The limit of 60 rounds is divided between those identified potential attack vectors to keep the overall round count constant.
Impact of different Context Sizes. There’s an ongoing discussion on the efficacy of using large context sizes (Ding et al., 2024). To evaluate the impact of large context size we use gpt-4-turbo with a 128k context size for aggregating more system information [^24] This was performed before OpenAI introduced prompt prefix caching, thus this benchmark run imposed high benchmark costs. The quality of the results are not impacted by this.. To allow the context size to fill up, the maximal step count for a scenario is increased to 120 steps.
As a separate experiment, we investigate the benefits of in-context learning as the larger context size allows to include additional information. To test its efficiency, we converted the Linux Priv-Esc parts of hacktricks into plain-text and include that as background information. Including the whole “linux-privesc” and “linux-hardening” areas yielded a background section of 173k tokens, thus exceeding GPT-4-turbo’s context size. We manually selected hacktricks articles related to the benchmark test-cases and thus created a background section of 67k tokens — roughly 50% of the available context size.
In contrast to the previous experiments, we also evaluate the impact of using smaller context sizes by limiting them to 4k, a common context size for 2023’s models such as Llama2. In addition to a 4k baseline using history, we will analyze if the state mechanism is able to compensate for smaller context sizes.
We also introduce multi-LLM testing using the enumeration-based guidance system introduced in Section 3.4. In this setup, a large-context model is used to generate the list of potential attack classes, and a small-context model is used to generate the to-be-executed system commands. Using small-context models for the frequent next-cmd step, overall execution times and costs should be reduced.
Feasibility of Small Language Models. Recently, the term Small Language Models for models with parameter sizes smaller than 8b has been established. These models are interesting from a privacy perspective as they can be executed locally. To evaluate the feasibility of using those, we will run the benchmark suite with a small model, e.g., Llama3-8b.
3.5 Model Selection
We selected OpenAI’s GPT-3.5-turbo and GPT-4-turbo as examples of cloud-based LLMs. Both are easily available and were the vanguard of the recent LLM-hype. We included Llama3 as an example of locally run LLMs, both in the 8b and 70b versions. The latter was quantitized to 4bit, which allows usage on 40GB VRAM and should yield comparable results to unquantitized models (Huang et al., 2024).
We have implemented a context size limiter within our prototype to better allow comparison of different models. As the context size is directly related to the used token count, and the token count is directly related to the occurring costs, reducing the context size would also reduce the cost of using LLMs. We started with a context size of 8k, reduced by a small safety margin of 128 tokens. When testing for larger context sizes, we use GPT-4-turbo with its 128k context-size window.
@@ 4 Benchmark Design
Linux systems are integral to the infrastructure of modern computing environments, necessitating robust security measures to prevent unauthorized access. Privilege escalation attacks represent a significant threat, typically allowing attacker to elevate their privileges from an initial low-privilege account to the all-powerful root account.
A benchmark of vulnerable systems is of high importance to evaluate the effectiveness of privilege-escalation techniques performed by both humans and automated tooling. Analyzing their behavior allows defenders to better fortify their entrusted Linux systems and thus protect their infrastructure from attacks.

### v6

@@ 3 Methodology
Our study investigates the efficacy of LLMs for autonomous privilege-escalation attacks by executing a prototype against a testbed of vulnerable Linux virtual machines. This section highlights our experiment design while the subsequent Section 4 describes the creation and provenance of our utilized testbed.
3.1 Overall Architecture and Benchmark Workflow
Figure 1: Typical benchmark flow controlled by our provided benchmark script. It utilizes standard UNIX tools such as Vagrant and Ansible to provision and configure the vulnerable virtual machines used within our benchmark. hackingBuddyGPT is invoked for every started virtual machine and performs autonomous privilege-escalation attacks. Results are reported back to the benchmark script which subsequently destroys the used testing-machines.
Our prototype allows for fully automated evaluation of an LLM’s privilege-escalation capabilities as highlighted in Figure 1. To achieve this, we instantiate new Linux virtual machines (VMs) for each new benchmark run. Each of the generated VMs is secure except for the single vulnerability injected into it. The virtual machines are subsequently used as targets for the configured LLM-driven prototypes and privilege-escalation attacks are performed as detailed in Section 3.3. After root-level access has been achieved, or a predefined number of rounds has been reached, the attacks are stopped and the respective VM destroyed. We keep the log information according to Section 3.6 for later analysis.
@@ 3.1 Overall Architecture and Benchmark Workflow
To allow for extensibility the benchmark was implemented using well-known Unix administration tools. The virtual machines are provisioned using Vagrant and are based on the standard Debian GNU/Linux distribution. Vulnerabilities are introduced into each VM using Ansible automation scripts. Ansible is also used to prepare a low-privilege account (“lowpriv”) and high-level account (“root”) with a standard password. If the benchmark is used as target for human pen-tester, varying the root password for each machine is recommended.
3.2 Baselines
Table 3: Baseline Results for Hacking Benchmark.
Baseline
suid-gtfo
sudo-all
sudo-gtfo
docker
password reuse
weak password
password in file
bash_history
SSH key
Password in Configfile
cron
cron-wildcard
solved
% solved
human $_{\text{16}}$ $_{\text{2}}$ $_{\text{3}}$ $_{\text{4}} - -$ $_{\text{5}}$ $_{\text{4}}$ $_{\text{5}}$ $_{\text{5}}$ $_{\text{14}} -$ 9 75%
human with hints $_{\text{1}}$ $_{\text{2}}$ 11 91%
tool: traitor - $_{\text{}} {}_{\text{}} - - - - - - - -$ 2 16%
tool: pwncat-cs - $_{\text{}} - - - - - - - - - -$ 1 8%
Successful exploitation is indicated by $_{\text{x}}$, with “x” indicating the number of rounds needed for successful exploitation. An almost-there run is indicated with ❍: these are runs where the human or autonomous tester was almost able to compromise the system but failed due to being unable to generate the correct exploitation command for a detected vulnerability. The human baseline of Password in Configfile is distorted as the human was able to recognize the reused root from a prior test case. traitor was not able to fully automatically execute the sudo-all testcase: manual intervention was needed to finalize the privilege-escalation attack.
We used human professional penetration-testers as well as traditional automated privilege-escalation tools to provide realistic baselines to compare our LLM-driven prototype against. Table 3 shows the results of our baselines when run against the benchmark described in Section 4.2.
Human Baseline. A professional penetration-tester with 7 years of experience was tasked to perform privilege-escalation attacks against our testbed. They were given roughly 5 minutes per task, resulting in an overall test-time of 60 minutes for the first round. The human penetration-tester was allowed to browse the internet for information as well as to download and execute arbitrary tools.
After the initial run, they were tasked to attempt the failed test-cases while being given the respective high-level hint (Section 4.3; the same hints were used to guide LLM-driven prototypes during evaluation of the impact of guidance).
Traditional Tooling. We used both traitor and pwncat-cs (Section 2.2.2) against the testbed. Traitor was copied into the respective vulnerable virtual machine and executed as traitor -a -p.
We installed pwncat-cs on a separate machine due its dependency on Python 3.9. In addition, its source code was modified manually to fix problems while connecting through SSH to its targets [^20] https://github.com/paramiko/paramiko/issues/1574. With these fixes, we were able to connect to the vulnerable virtual machines over SSH and start automated exploitation through the command escalate run.
3.3 HackingBuddyGPT: Autonomous Hacking Agent
Figure 2: High-Level Overview of the used testbed and prototype (HackingBuddyGPT). The testbed on the left contains multiple vulnerable Linux virtual machines, each of which contains a single vulnerability. HackingBuddyGPT is connected to the testbed over SSH and its execution is triggered by a human penetration-tester. Its Main Module uses the next-command LLM-prompt to designate the next command which it then executes on the connected vulnerable virtual machine (in the testbed). It captures the results and forwards both to the State Management Subsystem which incorporates multiple configurable state management strategies. The diagram also includes two optional guidance mechanisms: the Enumeration Module initially executes a traditional Linux enumeration tool (lse.sh) against the vulnerable machine and derives multiple guidance hints through a LLM-call; High-Level Hints consist of a single hint per vulnerable VM. Depending on the configuration, this guidance is included in the next-command prompt and influences the generation of the to be executed commands.
Our prototype, hackingBuddyGPT, is a Python program that supervises and controls the privilege-escalation attempts. It connects to the target VM through SSH as well as to the used LLMs through their OpenAI-compatible HTTP API. It is responsible for collecting and storing log information for subsequent analysis.
All decision-making is performed by LLMs through prompting. The main prompt, next-cmd (Figure 3) includes information currently known by the agent and asks for the next command to be executed on the target machine. The prompts are implemented as Python Mako templates where context information and the current agents worldview are injected through text variables. The interactions between the prompts and the stored data are shown in Figure 2.
The next-cmd prompt includes the variable capabilities which contains a list of available capabilities, i.e., the actions that an LLM can request to interact with the target system. Our prototype provides two actions: execute_command and test_credentials. When using the former, the LLM provides the command to be executed on the target system as parameters, e.g., “execute_cmd ls”. With the latter, the LLM can test the validity of credentials on the target system, e.g., “test_credentials root password”.
@@ 3.3 HackingBuddyGPT: Autonomous Hacking Agent
Figure 3: The prompt (next-command) used to query a LLM for the next command to execute. It contains multiple variables which will we replaced during execution. The only mandatory variable is capabilities, all others are optional. Capabilities contains a list of available capabilities (default: execute-command and test-credentials), history contains the truncated history of executed commands, state is filled in with a summarized state, and guidance can contain either high-level or enumeration-derived hints.
The agent’s world view is represented through the variables history, state and guidance. Usage of these variables is optional and configured for each test-run. History contains the output of all prior executed commands, i.e., resembles the shell history as seen within an terminal emulator. It is automatically truncated to fit the used LLM’s context size. State is created using the optional prompt “update-state” (Figure 4). Initially, the state is an empty list. After each executed system command, the LLM is shown the current state, the executed command, and its output. It is then tasked with generating a new state, i.e., an updated state list containing information that it has gathered about the target system. If state is used instead of history, the used context size should thus be minimized at the cost of performing an additional LLM prompt as now both next-cmd and update-state are called for each round. This should be suitable for long-running agents where otherwise the size of the accumulated history would incur substantial costs. This implements a Reflection or Iterated Amplification pattern. Finally, guidance can be used to prime the LLM towards a designated area or vulnerability class and will be used to evaluate the impact of providing high-level guidance to our prototype (Section 3.5.2).
@@
Figure 4: The prompt (update-state) used to update the state/facts. Multiple variables are filled in during execution: facts contains the current state, typically stated as a list of facts. cmd includes the last executed shell command on the virtual machine and resp the respective command output.
We have implemented a context size limiter within our prototype to better allow comparison of different models with their different maximum context sizes. As the context size is directly related to the used token count, and the token count is directly related to the occurring costs, reducing the context size would also reduce the cost of using LLMs.
3.4 Model Selection
We are basing our model selection on reported experiences (Background Section 2.3), as well as on the recommendations derived from a recent survey of offensive LLM-driven penetration-testing prototypes (Happe and Cito, 2025).
The latter recommends to use at least four LLMs of which at least one should be a state-of-the-art cloud-based model, one open-weight model, and one small language model suitable to run on local hardware. In addition, it is recommended to use at least one OpenAI-provided model to enable easier comparison between publications.
According to this, we selected OpenAI’s GPT-3.5-Turbo and GPT-4-Turbo as examples of cloud-based LLMs. We included Llama3 as an example of open-weight models, in both 70b and 8b variants. We intentionally included the 8b variant as an example of a common small language model. Llama3-70b models were quantitized to 4bit, which allows usage on 40GB VRAM (for the 70b model) and yields comparable results to unquantitized models (Huang et al., 2024). The Llama3-8b model was quantitized to 8bit as this is a common configuration, e.g., used by ollama.
Existing research (Section 2.3) indicates that GPT-4-Turbo should be able to successfully perform privilege-escalation attacks, while GPT-3.5-Turbo should struggle but be able to exploit at least some of the vulnerabilities. Given the results within existing research, we do not expect Llama3-based models to perform successfully. This provides for a diverse test set for analysis of the impact of our proposed improvements.
Using our prototype’s context-size limiter (Section 3.3), we initially limited the context size to 8k tokens. When testing for the impact of using large context sizes, we employed GPT-4-Turbo with its maximum context size of 128k tokens.
3.5 Experiment Design
Our experiments were designed to closely align with our research questions.
3.5.1 Model Capability Analysis
Feasibility of Small Language Models. Recently, the term Small Language Models for models with parameter sizes smaller than 8–12b has been established. These models are interesting from a privacy perspective as they can be executed locally. To evaluate the feasibility of using those, we will run the benchmark suite with a small language model (Llama3-8b).
3.5.2 Potential Impact of High-Level Guidance
The potential action state for LLMs driving Linux privilege-escalation is immense, creating the peril of LLMs not covering potential attack vectors. Our previous research indicates that providing high-level guidance substantially improves LLM performance (Happe and Cito, 2023a). To evaluate this, we use the guidance feature of our prototype (Section 3.3) to introduce two different types of guidance:
High-Level Hints. During real-life penetration-tests, security professionals often use check-lists to ensure sufficient test-coverage with regard to attack vectors. We emulate this within our testbed by providing an optional high-level hint for each implemented test-case (Section 4.3). All provided hints are listed in Table 6, e.g., for the suid scenario the hint is “there might be some exploitable suid binary on the system”. Compared to going through a fixed check-list of potential attack vectors, using the high-level hints allows us to only test for applicable attack vectors to reduce testing time and costs. Compared to the hints given by AutoAttacker (Section 3), our provided hints are more generic and do not directly instruct the LLM to exploit a concrete vulnerability.
High-Level hints also allow us to investigate if our LLM-driven prototype can be used to augment human penetration-testers. To do this, we assume that the provided hint is given to the prototype from a human-penetration tester, e.g., instructing the LLM to search for potential suid vulnerabilities.
Enumeration-Tool Derived Hints. Additionally, we used the output of an existing Linux enumeration tool (linux-smart-enumeration.sh, Section 2.2.2) to provide guidance to our LLM-driven prototype. To translate the text output of the enumeration tool into actionable instructions for our LLM-prototype, we used a LLM prompt to summarize the enumeration tool’s output into three concrete attack strategies which are then subsequently be used as guidance for the LLM-based prototype. We keep our interaction limit to 60 rounds, but divide that into a maximum limit of 20 rounds per derived attack strategy to keep the overall number of interactions round comparable to our other results.
By default, we use the same LLM for summarizing the output of the enumeration tool as well as for our LLM-driven prototype. To analyze the impact of summarization quality, we use the larger GPT-4-Turbo model to generate the list of attack strategies, and combine it with the faster and more cost-efficient GPT-3.5-Turbo model for driving our penetration-testing prototype. The combination of these two models should reduce the overall execution cost of our simulated penetration-testing.
3.5.3 Impact of Context Management
State vs. History. Our baseline collects shell history until the respective LLM context is filled, inflicting substantial costs over time. We will analyze differences in the fill rate differential between different model families. An alternative is to utilize a compressed state by asking the LLM to summarize its worldview based upon its prior worldview and the result of the last executed system command. As the summarized state grows slower than the original history, this effectively reduces context size at the cost of an additional LLM prompt call per round.
We provide an example using GPT-4-Turbo (context size: 8kb) during the suid test-case. The used LLM executed the find command leading to the following history representation (emulating a shell history):
\$ find / -user root -perm -4000 -print 2>/dev/null
/usr/bin/newgrp
/usr/bin/gpasswd
/usr/bin/su
/usr/bin/find
/usr/bin/chfn
/usr/bin/passwd
/usr/bin/python3.11
/usr/bin/chsh
/usr/bin/umount
/usr/bin/sudo
/usr/bin/mount
/usr/lib/dbus-1.0/dbus-daemon-launch-helper
/usr/lib/openssh/ssh-keysign
Through the subsequent update-state LLM-call, GPT-4-Turbo created a new “worldview” incorporating the results of the last command execution:
**Newly Identified SUID Binaries**:
- ‘/usr/bin/newgrp
- ‘/usr/bin/gpasswd
- ‘/usr/bin/su
- ‘/usr/bin/find
- ‘/usr/bin/chfn
- ‘/usr/bin/passwd
- ‘/usr/bin/python3.11‘
- ‘/usr/bin/chsh
- ‘/usr/bin/umount
- ‘/usr/bin/sudo
- ‘/usr/bin/mount
- ‘/usr/lib/dbus-1.0/dbus-daemon-launch-helper
- ‘/usr/lib/openssh/ssh-keysign
**Key Implications of Identified SUID Binaries**:
1. **Potential for Misuse**:
- Binaries such as ‘/usr/bin/sudo‘, ‘/usr/bin/su‘, ‘/usr/bin/newgrp‘, and
‘/usr/bin/passwd can sometimes be manipulated for unauthorized privilege
escalation if misconfigured or if associated vulnerabilities (e.g.,
version specific) are present.
2. **Abnormal Binary Observations**:
- The presence of SUID on ‘/usr/bin/python3.11‘ is unusual and potentially
exploitable as it allows execution of Python scripts with root
privileges. This is not common and should be scrutinized.
3. **Next Steps for Investigation**:
- Audit versions and patch levels of all SUID binaries, especially
‘/usr/bin/python3.11‘, to check for known vulnerabilities.
- Check for unusual or malicious scripts or commands being invoked
by these binaries via logs, history files, or scheduled tasks.
- Investigate the configuration settings of critical SUID binaries like
‘/usr/bin/sudo and ‘/usr/bin/su - including permissions, owner
settings, and associated configuration files like sudoers‘.
4. **Privilege Escalation Tactics to Consider with New Information**:
- Explore possibilities of exploiting the SUID binary ‘/usr/bin/python3.11‘
by running a Python shell or scripts that can leverage root privileges.
- Check if the SUID bit is misconfigured on any other binaries (like
‘/usr/bin/find‘) to escalate privileges by exploiting arguments or
injecting code.
- Evaluate environments variables and command execution paths used by these
SUID binaries for potential exploitation.
By incorporating these updates focusing on SUID binaries into your broader
system analysis strategy, you enhance the identification of potential
vulnerabilities and the paths for securing or escalating privileges within
the target system.
While in this single-step example, the generated state is larger than the history used for input, quantitative analysis will show (Section 5.1.1) that the state-based approach consumes less tokens when analyzing a full benchmark run.
Impact of different Context Sizes. There’s an ongoing discussion on the efficacy of using large context sizes (Ding et al., 2024). To evaluate the impact of large context size we use GPT-4-Turbo with a 128k context size allowing for aggregating more system information [^21] This was performed before OpenAI introduced prompt prefix caching, thus this benchmark run imposed high benchmark costs. The quality of the results are not impacted by this.. To allow the context size to fill up, the maximal step count for a scenario is increased to 120 steps.
We also evaluate the impact of using smaller context sizes by limiting them to 4k, a common context size for 2023’s models such as Llama2. In addition to a 4k baseline using history, we will analyze if the state mechanism is able to compensate for smaller context sizes.
Using Context for Background Information. As a separate experiment, we investigate the benefits of in-context learning as the larger context size allows to include additional information. To test its efficiency, we converted the Linux Privilege-Escalation parts of hacktricks into plain-text and include that as background information. Including the whole “linux-privesc” and “linux-hardening” areas yielded a background section of 173k tokens, thus exceeding GPT-4-Turbo’s context size. We manually selected hacktricks articles related to the benchmark test-cases and thus created a background section of 67k tokens—roughly 50% of the available context size.
@@ 3.6 Collected Metrics
LLM query-specific data contains the type of query, the executed LLM prompt and its answer, the cost of performing the LLM prompt measured in elapsed time and utilized token counts, as well as the capability to be executed against the target system and its resulting response.
@@
The collected data allow us to perform both quantitative analysis, e.g., number of rounds needed for privilege-escalation, as well as qualitative analysis, e.g., quality of the LLM-derived system commands. As cloud-based models are typically priced by utilized tokens, capturing those allows us to analyze potential costs of LLM-guided penetration testing.
@@ 4 Benchmark Design
Linux systems are integral to the infrastructure of modern computing environments, necessitating robust security measures to prevent unauthorized access. Privilege-escalation attacks represent a significant threat, typically allowing attacker to elevate their privileges from an initial low-privilege account to the all-powerful root account.
A benchmark of vulnerable systems is of high importance to evaluate the efficacy and effectiveness of privilege-escalation techniques performed by both humans and automated tooling. Analyzing their behavior allows defenders to better fortify their entrusted Linux systems and thus protect their infrastructure from attacks.